OpenStudy (anonymous):

solve for t t+1=x(2^t(t-1)t)

OpenStudy (anonymous):

\[t+1=x(2^t(t-1)t) \] seems like it?

OpenStudy (anonymous):

θ(1+x)^-(1+θ),x>1

OpenStudy (anonymous):

can someone intergrate from 1to infinity

OpenStudy (anonymous):

t+1=x(2^t(t-1)t) t =x(2^(t(t)-t)(t)) is it like ths

OpenStudy (anonymous):

\[t+1=x(2t ^{2}(t-1))\]

OpenStudy (experimentx):

it's a cubic equation in t.

OpenStudy (anonymous):

yes but i don't get t

OpenStudy (experimentx):

there is a brute method http://en.wikipedia.org/wiki/Cubic_function#Cardano.27s_method

OpenStudy (anonymous):

wen i simplify i got \[\frac{ (t+1) }{ t-1 }=x(2t ^{2})\]

OpenStudy (experimentx):

this doesn't look easy http://www.wolframalpha.com/input/?i=solve++t%2B1%3Dx%28t^2%28t-1%29%29+for+t

OpenStudy (experimentx):

\[ 2x t^3 - 2x t^2 - t - 1 = 0\] directly use this http://en.wikipedia.org/wiki/Cubic_function#General_formula_of_roots to get the roots of cubic equation. or, make this type of substitution http://en.wikipedia.org/wiki/Cubic_function#Reduction_to_a_depressed_cubic to get cubic in depressed form and use Cardano's method to get the roots http://en.wikipedia.org/wiki/Cubic_function#Cardano.27s_method

OpenStudy (experimentx):

this might be easy for particular value of x.

OpenStudy (anonymous):

I suppose there is something wrong somewhere. I'm thinking the t + 1 is really t - 1. (that is my thinking)